OpenStudy (anonymous):

L = 3in w= 5 in and 3in find w?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[5/3=3/x\] Cross multiply to get \[5x = 9\] solve for x to get 9/5

OpenStudy (anonymous):

5X 3 = 15 = 3x = 15/ 3 = x = 5

OpenStudy (anonymous):

?? is that right?

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

No, it would look like this: 5/3=3/x you first multiply the 5 on the top left with the x on the bottom right. then multiply the 3 on the bottom left with the 3 on the bottom right

OpenStudy (anonymous):

then you would have 5x=9 can you solve from there?

OpenStudy (anonymous):

i got 5x =9

OpenStudy (anonymous):

Yes, now divide both sides by 5... and what's your answer?

OpenStudy (anonymous):

how can u divde 9/5 ? i got 1.8?

OpenStudy (anonymous):

or you can simply leave it as an improper fraction.. whatever your teacher says to do. and yes it would be 1.8 if needed

OpenStudy (anonymous):

so its 14/5 ?